# Multi-Model Framework for Reconstructing Gamma-Ray Burst Light Curves

### v1

@@ 2.3 Dataset Preprocessing
We follow the same preprocessing and training pipeline as described in Dainotti et al. 2023c; Manchanda et al. 2024, which ensures consistency across all models and facilitates a fair comparison. We used a random subset of 16 GRBs, four from each GRB category, to tune our model by minimizing the Mean Squared Error (MSE) loss. This tuning was conducted using optuna framework Akiba et al. (2019) and the parameters obtained were used for the reconstruction of the entire 545 GRBs dataset. After obtaining the parameters, each model is trained independently on individual GRB LCs, using the log10-transformed time and flux values. The data is scaled to [0,1] using min-max normalization, described in Eq. 2:
@@
where $X_{min}$ and $X_{max}$ are the minimum and the maximum value in $D$.
After training the model and reconstructing the temporal gaps, we estimate the aleatoric uncertainty in the reconstructed flux values in a similar way as Dainotti et al. 2023c; Manchanda et al. 2024. For each model, we perform 1000 Monte Carlo simulations by adding noise sampled from a Gaussian distribution, fitted to the residuals between the original and reconstructed fluxes. This captures the inherent data variability of the LCs. The 95% confidence interval (CI) is computed using the 2.5th and 97.5th percentiles of the distribution of the reconstructed values at each time point, providing a reliable estimate of the aleatoric uncertainty.
@@ 2.4.1 Deep Gaussian Model
This method layers the GP (Rasmussen and Williams, 2006) regression technique, enabling deep learning for probabilistic outputs tailored to regression tasks. A Deep Gaussian Process (DGP) Damianou and Lawrence 2013 extends the traditional GP framework by stacking multiple GP layers, where each layer models latent functions whose outputs serve as inputs to the next. Unlike a single-layer GP, which assumes a joint Gaussian distribution between inputs and outputs, a DGP introduces intermediate layers of latent variables—unobserved, inferred quantities that capture complex underlying structure—breaking the joint Gaussianity. This hierarchical setup enables the model to capture increasingly complex patterns in LC data by learning a distribution over the latent space at each layer. As shown in Fig. 2, we can view the DGP as a form of Markov Chain, where each layer conditionally depends only on its immediate predecessor, thus simplifying the propagation of uncertainty through the network.
@@
In our model, we use two DGP layers, each consisting of a scalable single-layer approximate GP. The model is implemented using GPyTorch, a Gaussian Process library built on top of PyTorch. The input is standardized using Eq. 2 before training to ensure stable optimization. For smooth and effective transitions, we use the Radial Basis Function (RBF) kernel with a length scale of 2.0. The RBF kernel, also known as the squared exponential kernel, measures similarity between inputs based on their Euclidean distance and assumes that points closer in the input space are more strongly correlated. Within a single-layer GP, we employ a GPyTorch-enabled Cholesky variational strategy to learn the optimal inducing point locations. Inducing points are a small set of representative inputs that help approximate the full GP, making it faster and more scalable. The Cholesky strategy improves stability and efficiency by factorizing the covariance matrix, allowing the model to handle larger datasets without sacrificing performance. We use 50 inducing points in the first layer and 30 for the second layer, selected through manual hyperparameter tuning on different GRB types. We tested several combinations on all types of GRBs and selected the configuration that minimized reconstruction error while maintaining computational efficiency. All evaluations were done under consistent training settings and preprocessing steps to ensure a fair comparison. Each layer outputs a latent distribution whose mean is passed as input to the next layer, modeling a two-layered function:
@@
where $\mathbf{y}$ denotes the observed outputs, $\mathbf{f}$ represents the latent function outputs, $\sigma_{n}^{2}$ is the learnable noise variance, and $\mathbf{I}$ is the identity matrix, assuming Gaussian noise. The noise is initialized at 0.01 and constrained within the interval $[10^{-4},0.05]$ to prevent the model from overfitting or underfitting the noise levels across different GRB types.
@@ 2.4.2 Temporal Convolutional Neural Network
The Temporal Convolutional Network (TCN) (Bai et al., 2018b) is a convolutional neural network architecture designed for sequence modeling tasks. TCN uses causal convolutions, where the output at time step t is computed exclusively from elements at time t and earlier, preventing information leakage from future time steps and enhancing convolutions through dilation, where filters (learnable weights that detect patterns) are applied with regularly spaced gaps across the input sequence, as illustrated in Fig. 3(b). For a 1D input sequence $\mathbf{x}\in\mathbb{R}^{n}$ and a filter $f:\{0,\dots,k-1\}\rightarrow\mathbb{R}$, the dilated convolution $F$ at position $s$ is given by:
@@
Residual connections (He et al., 2016) involve adding the input of a layer directly to its output, skipping one or more intermediate layers as shown in Fig. 3(a). This technique allows the model to learn identity mappings, which helps prevent the vanishing gradient problem and facilitates more stable gradient flow during backpropagation. The TCN uses residual connections to improve training stability and ensure efficient convergence.
Unlike RNNs, which process sequences step by step, TCNs operate over the entire sequence in parallel. This parallelism improves computational efficiency and reduces training time. In terms of resource efficiency, TCNs generally require less memory during training compared to gated RNNs such as LSTMs (Hochreiter and Schmidhuber, 1997b) and GRUs (Cho et al., 2014). This is because TCNs share convolutional filters across time steps and do not need to store internal states or gate activations.
In this study, the TCN was implemented using the keras-tcn library in combination with Keras’ Functional API. The network was created with 64 filters, a kernel size of 5, a dropout rate of 0.1, a single stack of residual blocks, and dilation factors set to $(1,2,4,8,16,32)$. The TCN output was passed to a Dense layer with a linear activation function. The input shape was $(1,1,1)$.
Hyperparameter optimization was performed through grid search on a randomly selected subset of GRB data, using the following ranges:
@@
The ideal set of values for all GRBs was determined by combining the values that performed the best on this subset. Using the Adam optimizer with a batch size of 8 and MSE as the loss function, the model was trained across 100 epochs. Both input features and target values were normalized using MinMaxScaler to ensure consistent scaling.
Figure 3: (a) Residual block of a Temporal Convolutional Network. (b) Expanded view of a dilated convolution layer
2.4.3 Convolutional Neural Network - Long Short Term Memory
Convolutional Neural Networks (CNNs) (LeCun et al., 1998; Zhang et al., 2019) and Bidirectional Long Short-Term Memory (Bi-LSTM) networks have individually demonstrated strong performance in extracting spatial and temporal features, respectively, making their hybrid combination particularly powerful for sequential pattern learning. CNN-BiLSTM models leverage this synergy by integrating convolutional feature extraction with temporal sequence modeling, enabling superior performance in tasks involving complex and noisy time-series data (Wang et al., 2017; Yao et al., 2017).
In the CNN-BiLSTM architecture, the input LC is initially passed through one or more convolutional layers that act as localized feature detectors. These layers apply multiple filters to capture short-range dependencies and local patterns such as flares, peaks, or breaks in GRB LCs. Filters are like pattern detectors that slide over the input data and look for patterns in a local region of the input. This local processing is effective in denoising the signal and emphasizing key morphological structures relevant for downstream reconstruction (Kiranyaz et al., 2021). A typical convolution operation in a 1D CNN can be defined as:
@@ 2.4.3 Convolutional Neural Network - Long Short Term Memory
The compressed feature maps from the CNN are then fed into a Bi-LSTM block that processes the sequence in both forward and backward directions. This dual perspective allows the model to capture long-range dependencies while incorporating context from both past and future time steps (Graves et al., 2005). For GRB LCs, where information from both earlier and later intervals can improve reconstruction accuracy, especially in the presence of complex events like multiple flares, this bidirectional processing is critical.
@@
    A 1D convolutional layer with 64 filters, kernel size $k=3$, ReLU activation, He-uniform initialization, and $L_{2}$-regularization.
@@
    Three stacked Bi-LSTM layers (100 units each).
@@
We use the Adam optimizer with a learning rate of 0.0001 for training. This hyperparameter was determined during the hyperparameter tuning phase as mentioned in the section 2.3.
@@ 2.4.4 Bayesian Neural Network
This converts the inference problem into an optimization one, often using reparameterization tricks to enable backpropagation through stochastic variables. Dropout-based approximations, such as Monte Carlo Dropout, implement a form of variational inference where dropout is applied at test time, sampling from an implicit posterior over the weights. These samples generate predictive distributions from which uncertainty can be estimated.
@@
The model is implemented using PyTorch and torchbnn, structured with hidden layers (num_layers). Each hidden layer is a BNNHiddenLayer, which wraps a BNNLinear layer with trainable Gaussian-distributed weights and biases. The hidden layers are defined by an activation function (Leaky ReLU, Tanh, or Swish) and include dropout for regularization. The output consists of two parallel linear heads - one for the predictive and another for the log-variance. Hyperparameters such as the number of layers, hidden units, dropout rate, and activation function are optimized using Optuna. The values of PRIOR_MU and PRIOR_SIGMA are set to 0.0 and 0.1 respectively. The training process involves forward propagation with stochastic weight sampling, minimizing a composite loss function consisting of:
@@
$\mathcal{L}(\theta)=\underbrace{-\left[\log p(D\mid\theta)\right]}_{\text{Negative Log-Likelihood}}+\underbrace{\mathrm{KL}\left(q(\theta)\,\|\,p(\theta)\right)}_{\text{KL Divergence}}.$ (12)
The KL-divergence term is scaled by a fixed multiplicative factor $\lambda_{\text{KL}}=10^{-3}$ to balance model fit and uncertainty. After hyperparameter optimization, the model is retrained on the full LC of every GRB separately. For prediction, multiple stochastic forward passes are conducted to generate a distribution of outputs, from which prediction intervals and uncertainty estimates are derived.
@@ 2.5 The statistical approach
This section provides a brief overview of various statistical-based models for LC reconstruction
@@ 2.5.1 Polynomial Curve Fitting
In previous sections, we experimented with various sophisticated approaches. However, now, we aim to simplify the curve-fitting process. Our approach involves fitting an n-degree polynomial to the LC, starting with an initial fine-tuning phase. Similar polynomial regression techniques have been employed in previous studies, as noted by (AlShammari, 2024). An n-degree polynomial can be defined as:
@@
The objective is to select the degree $n$ such that the model achieves an optimal balance between flexibility and generalization. To do this, we apply MinMaxScaling given by the Eq. 2 and then randomly sample four instances from each GRB class and determine the polynomial degree that minimizes MSE loss while maintaining model simplicity. In our experiments, a third-degree polynomial yielded the best performance. Apart from choosing the polynomial degree, we also fine-tune the initial parameters $c1$, $c2$, and $c3$, which serve as the starting point for optimization. The resulting polynomial equation takes the form:
$y=-0.354+0.796x-0.0255x^{2}.$ (14)
This polynomial is then fitted to the data using the $scipy.optimize.curve\_fit$ function, which estimates the optimal coefficients $\{c1,c2,c3\}$ for each GRB.
@@ 2.5.2 Isotonic regression
This approach uses a statistical, non-parametric method to estimate monotonic (either decreasing or increasing) one-dimensional data. Unlike classical regression which assumes a linear functional form, the non-parametric approach does not presume the structure. Therefore, this method is more flexible than standard regression and fits the data more accurately as it is displayed in Fig. 7.
Given that the LCs generally show a decreasing trend over time, in our implementation, a non-increasing isotonic regression model is applied (Best and Chakravarti, 1990).
@@
$x_{1}<x_{2}<\ldots<x_{n}$, and for any $i>j$, it holds that $y_{j}\geq y_{i}$. With such assumption, the isotonic regression seeks a non-increasing function $f:X\rightarrow\mathbb{R}$ that minimizes the weighted least squares criterion:
@@
where $w_{j}\geq 0$ is added weight to the observation.
Figure 7: Weighted least-square method minimizes the residuals which are marked with blue lines
@@
To find the most optimal function $f$ the pool adjacent violators (PAV) algorithm is used. This procedure is implemented in Python sklearn package in IsotonicRegression in the fitting method (de Leeuw et al., 2009). Let us present this algorithm briefly.
@@
Therefore the sequence $\hat{y}_{1},\ldots,\hat{y}_{n}$ are the fitted values resulting from isotonic regression, such that they satisfy the monotonicity $\hat{y}_{1}\geq\hat{y}_{2}\geq\ldots\geq\hat{y}_{n}$.
After training the model, the procedure of predicting new value $x$ is given by:
@@
3 Results
Decreasing the uncertainty associated with the model parameters is an objective of LCR. To evaluate this, the error fractions, symbolized by $EF$, are calculated for every model parameter in both the primary datasets and after reconstructing them using the model. Eq. 22, 23, and 24 reflect the error fractions for the three parameters as stated in Dainotti et al. 2023c:

### v2

@@ 2.3 Dataset Preprocessing
We follow a similar preprocessing and training pipeline as described in dainotti2023stochastic; manchanda2025gammarayburstlightcurve, which ensures consistency across all models and facilitates a fair comparison. We used a random subset of 16 GRBs, four from each GRB category, to tune our model by minimizing the Mean Squared Error (MSE) loss. This tuning was conducted using optuna framework (2019optuna) and the parameters obtained were used for the reconstruction of the entire 521 GRBs dataset. After obtaining the parameters, models, which required automatic hyperparameter tuning, are trained independently on individual GRB LCs, using the log10-transformed time and flux values. The data is scaled to [0,1] using min-max normalization, described in Eq. 2:
@@
where $X_{min}$ and $X_{max}$ are the minimum and the maximum value in the training dataset $D$.
For QSS & DGP Z-score normalization was performed, described as:
$X_{i}=\frac{X_{i}-X_{mean}}{X_{std}};i\in D,$ (3)
where $X_{mean}$ and $X_{std}$ are the mean and standard deviation of the training dataset $D$.
After training the model and reconstructing the temporal gaps, we estimate the aleatoric uncertainty in the reconstructed flux values in a similar way as dainotti2023stochastic; manchanda2025gammarayburstlightcurve. DGP and QSS models do not require Monte Carlo (MC) simulations for uncertainty quantification, while for the BNN and TCN models, 100 MC simulations were done, while for the rest of the models, 1000 MC simulations were made. This captures the inherent data variability of the LCs. The 95% confidence interval (CI) is computed using the 2.5th and 97.5th percentiles of the distribution of the reconstructed values at each time point, providing a reliable estimate of the aleatoric uncertainty. In CNN Bi-LSTM, Polynomial Curve Fitting, TCN, and Isotonic Regression models, uncertainty are computed from the observational errors of input data. Deep GP and QSS models’ uncertainty is calculated from the residuals.
@@ 2.4.1 Deep Gaussian Model
This method layers the GP (rasmussen2006gp) regression technique, enabling deep learning for probabilistic outputs tailored to regression tasks. A Deep Gaussian Process (DGP) (damianou2013deep) extends the traditional GP framework by stacking multiple GP layers, where each layer models latent functions whose outputs serve as inputs to the next. Unlike a single-layer GP, which assumes a joint Gaussian distribution between inputs and outputs, a DGP introduces intermediate layers of latent variables—unobserved, inferred quantities that capture complex underlying structure—breaking the joint Gaussianity. This hierarchical setup enables the model to capture increasingly complex patterns in LC data by learning a distribution over the latent space at each layer. As shown in Fig. 2, we can view the DGP as a form of Markov Chain, where each layer conditionally depends only on its immediate predecessor, thus simplifying the propagation of uncertainty through the network.
@@
In our model, we use two DGP layers, each consisting of a scalable single-layer approximate GP. The model is implemented using GPyTorch, a Gaussian Process library built on top of PyTorch. The input is standardized before training to ensure stable optimization. For smooth and effective transitions, we use the Radial Basis Function (RBF) kernel with a length scale of 2.0. The RBF kernel, also known as the squared exponential kernel, measures similarity between inputs based on their Euclidean distance and assumes that points closer in the input space are more strongly correlated. Within a single-layer GP, we employ a GPyTorch-enabled Cholesky variational strategy to learn the optimal inducing point locations. Inducing points are a small set of representative inputs that help approximate the full GP, making it faster and more scalable. The Cholesky strategy improves stability and efficiency by factorizing the covariance matrix, allowing the model to handle larger datasets without sacrificing performance. We use 35 inducing points in the first layer and 25 for the second layer, selected through manual hyperparameter tuning on different GRB types. We tested several combinations on all types of GRBs and selected the configuration that minimized reconstruction error while maintaining computational efficiency. All evaluations were done under consistent training settings and preprocessing steps to ensure a generalization comparison. Each layer outputs a latent distribution whose mean is passed as input to the next layer, modeling a two-layered function:
@@
where $\mathbf{y}$ denotes the observed outputs, $\mathbf{f}$ represents the latent function outputs, $\sigma_{n}^{2}$ is the learnable noise variance, and $\mathbf{I}$ is the identity matrix, assuming Gaussian noise. The noise is initialized at 0.01 and constrained within the interval $10^{-4},0.05]$ to prevent the model from overfitting or underfitting the noise levels across different GRB types.
@@ 2.4.2 Temporal Convolutional Neural Network
The Temporal Convolutional Network (TCN) (Bai2018AnEE) is a convolutional neural network architecture designed for sequence modeling tasks. TCN uses causal convolutions, where the output at time step t is computed exclusively from elements at time t and earlier, preventing information leakage from future time steps and enhancing convolutions through dilation, where filters (learnable weights that detect patterns) are applied with regularly spaced gaps across the input sequence. Residual connections (he2016deep) involve adding the input of a layer directly to its output, skipping one or more intermediate layers as shown in Fig. 3(a). The expanded view of the dilation convolution layer is illustrated in Fig. 3(b). For a 1D input sequence $\mathbf{x}\in\mathbb{R}^{n}$ and a filter $f:\{0,\dots,k-1\}\rightarrow\mathbb{R}$, the dilated convolution $F$ at position $s$ is given by:
@@
This technique allows the model to learn identity mappings, which helps prevent the vanishing gradient problem and facilitates more stable gradient flow during backpropagation. The TCN uses residual connections to improve training stability and ensure efficient convergence.
Unlike RNNs, which process sequences step by step, TCNs operate over the entire sequence in parallel. This parallelism improves computational efficiency and reduces training time. In terms of resource efficiency, TCNs generally require less memory during training compared to gated RNNs such as LSTMs (Hochreiter1997LongSM) and Gate Recurrent Unit, GRUs, (Cho2014LearningPR). This is because TCNs share convolutional filters across time steps and do not need to store internal states or gate activations.
In this study, the TCN was implemented using the keras-tcn library in combination with Keras’ Sequential. The network was created with 64 filters, a kernel size of 5, a dropout rate of 0.1, a single stack of residual blocks, and dilation factors set to $(1,2,4,8,16,32)$. The TCN output was passed to a Dense layer with a linear activation function of shape $(1,1,1)$ where the first, second and third dimensions represent batch size, the time stamps, and the input, respectively. Hyperparameter optimization was performed through grid search on a randomly selected subset of GRB data, using the following ranges:
@@
The ideal set of values for all GRBs was determined by combining the values that performed the best hyperparameters on this subset. Using the Adam optimizer with a batch size of 8 and MSE as the loss function, the model was trained across 100 epochs. Both input features and target values were normalized using MinMaxScaler to ensure consistent scaling.
Figure 3: a) Residual block of a Temporal Convolutional Network. (b) Expanded view of a dilated convolution layer
2.4.3 Convolutional Neural Network - Bidirectional Long Short Term Memory
Convolutional Neural Networks (CNNs) (lecun1998gradient; zhang2019cnn) and Bidirectional Long Short-Term Memory (Bi-LSTM) networks have individually demonstrated strong performance in extracting spatial and temporal features, respectively, making their hybrid combination particularly powerful for sequential pattern learning. CNN-BiLSTM models leverage this synergy by integrating convolutional feature extraction with temporal sequence modeling, enabling superior performance in tasks involving complex and noisy time-series data (wang2017time; yao2017deepsense).
In the CNN-BiLSTM architecture, the input LC is initially passed through one or more convolutional layers that act as localized feature detectors. These layers apply multiple filters to capture short-range dependencies and local patterns such as flares, peaks, or breaks in GRB LCs. Filters are like pattern detectors that slide over the input data and look for patterns in a local region of the input. This local processing is effective in denoising the signal and emphasizing key morphological structures relevant for downstream reconstruction (kiranyaz20211dcnn). A typical convolution operation in a 1D CNN can be defined as:
@@ 2.4.3 Convolutional Neural Network - Bidirectional Long Short Term Memory
The compressed feature maps from the CNN are then fed into a Bi-LSTM block that processes the sequence in both forward and backward directions. This dual perspective allows the model to capture long-range dependencies while incorporating context from both past and future time steps (graves2005bidirectional). For GRB LCs, where information from both earlier and later intervals can improve reconstruction accuracy, especially in the presence of complex events like multiple flares, this bidirectional processing is critical.
@@
    A 1D convolutional layer with 64 filters, kernel size $k=3$, ReLU activation, He-normal initialization, and $L_{2}$-regularization.
@@
    Three stacked Bi-LSTM layers (100 units each, ”100 units” refers to the number of hidden units (or neurons) within that specific LSTM layer.), utilizing ReLU activation for the first layer and Swish for the subsequent two. =The ReLU activation introduces non-linearity while being computationally efficient, enabling the model to capture temporal patterns without suffering from vanishing gradient issues. The ReLU function outputs zero for negative inputs and returns the input value for positive ones. In contrast, the Swish activation, defined as $f(x)=x\cdot sigmoid(x)$, often performs better in deeper networks because it is smoother and permits small negative activations instead of truncating them completely. The sigmoid component squashes its output values to the range (0, 1), thereby modulating the gradient flow more smoothly during training.
@@
We use the Adam optimizer with a learning rate of 0.0001 for training. This hyperparameter was determined during the hyperparameter tuning phase as mentioned in section 2.3.
@@ 2.4.4 Bayesian Neural Network
This converts the inference problem into an optimization one, often using reparameterization tricks to enable backpropagation through stochastic variables. Dropout-based approximations, such as Monte Carlo Dropout, implement a form of variational inference where dropout is applied at test time, sampling from an implicit posterior over the weights. These samples generate predictive distributions from which uncertainty can be estimated. A detailed architecture is illustrated in Figure 5.
@@
The model is implemented using PyTorch and torchbnn, structured with hidden layers (num_layers). The first layer is a linear layer for mapping the input dimension to the subsequent hidden layers. Each layer afterwards is a BNNHiddenLayer, which wraps a BNNLinear layer with trainable Gaussian-distributed weights and biases. The hidden layers are defined by an activation function (Leaky ReLU, Tanh, or Swish) and include dropout for regularization. The output consists of a single Bayesian linear layer with two output features, which are subsequently interpreted as the predictive mean and the log-variance. Hyperparameters such as the number of layers, hidden units, dropout rate, and activation function are optimized using Optuna, see Table 1. The values of PRIOR_MU and PRIOR_SIGMA are set to 0.0 and 0.1, respectively. The training process involves forward propagation with stochastic weight sampling, minimizing a composite loss function consisting of:
@@
  • 3.
    Weighted MSE loss for aligning the predicted value ($\hat{y}$) with true value ($y$) according to the weights ($w_{i}$) assigned for each predicted class ($i$).
  • 4.
    Additionally, a log-variance penalty term is added to prevent the predicted variances from becoming excessively large or unstable during training.
$\mathcal{L^{\prime}}(\theta)=\underbrace{-\left[\log p(D\mid\theta)\right]}_{\text{Negative Log-Likelihood}}+\underbrace{\mathrm{KL}\left(q(\theta)\,\|\,p(\theta)\right)}_{\text{KL Divergence}}+\underbrace{\frac{\sum_{i=0}^{n}{w_{i}(y-\hat{y})^{2}}}{\sum_{i=0}^{n}w_{i}}}_{\text{Weighted MSE Loss}}$ (13)
$\mathcal{L}(\theta)=\mathcal{L^{\prime}}(\theta)+\underbrace{\lambda_{\text{log-var}}\sum_{i=1}^{N}{(\log{\sigma^{2}}})^{2}}_{\text{Log-Variance Penalty}}.$ (14)
The KL-divergence term is scaled by a fixed multiplicative factor $\lambda_{\text{KL}}=10^{-3}$ to balance model fit and uncertainty. After hyperparameter optimization, the model is retrained on the full LC of every GRB separately. For prediction, multiple stochastic forward passes are conducted to generate a distribution of outputs, from which prediction intervals and aleatoric uncertainty estimates are derived.
@@ 2.5 The statistical approach
This section provides a brief overview of various statistical-based models for LC reconstruction.
@@ 2.5.1 Polynomial Curve Fitting
In previous sections, we analyzed with various sophisticated approaches. However, now, we aim to simplify the curve-fitting process. Our approach involves fitting an n-degree polynomial to the LC, starting with an initial fine-tuning phase. Similar polynomial regression techniques have been employed in previous studies, as noted by (10.5120/ijca2024923400). An n-degree polynomial can be defined as:
@@
The objective is to select the degree $n$ such that the model achieves an optimal balance between flexibility and generalization. To do this, we apply MinMaxScaling given by the Eq. 2 and then randomly sample four instances from each GRB class and determine the polynomial degree that minimizes MSE loss while maintaining model simplicity. In our analysis, a second-degree polynomial yielded the best performance. Apart from choosing the polynomial degree, we also fine-tune the initial parameters $c1$, $c2$, and $c3$, which serve as the starting point for optimization. The resulting polynomial equation takes the form:
$y=-0.35+0.80x-0.026x^{2}.$ (16)
This polynomial is then fitted to the data using the scipy.optimize.curve_fitscipy.optimize.curve\_fit function, which estimates the optimal coefficients $\{c1,c2,c3\}$ for each GRB.
@@ 2.5.2 Isotonic regression
This approach uses a statistical, non-parametric method to estimate monotonic (either decreasing or increasing) one-dimensional data. Unlike classical regression, which assumes a linear functional form, the non-parametric approach does not presume the structure. Therefore, this method is more flexible than standard regression and fits the data more accurately, as it is displayed in Fig. 7.
Given that the LCs generally show a decreasing trend over time, in our implementation, a non-increasing isotonic regression model is applied (Best1990).
@@
$x_{1}<x_{2}<\ldots<x_{n}$, and for any $i>j$, it holds that $y_{j}\geq y_{i}$. With such an assumption, the isotonic regression seeks a non-increasing function $f:X\rightarrow\mathbb{R}$ that minimizes the weighted least squares criterion:
@@
where $w_{j}\geq 0$ is the added weight to the observation.
Figure 7: Weighted least-square method minimizes the residuals, which are marked with blue lines
@@
To find the most optimal function $f$, the pool adjacent violators (PAV) algorithm is used. This procedure is implemented in the Python sklearn package in IsotonicRegression in the fitting method (deLeeuw2009). Let us present this algorithm briefly.
@@
Therefore, the sequence $\hat{y}_{1},\ldots,\hat{y}_{n}$ is the fitted values resulting from isotonic regression, such that they satisfy the monotonicity $\hat{y}_{1}\geq\hat{y}_{2}\geq\ldots\geq\hat{y}_{n}$.
After training the model, the procedure of predicting a new value $x$ is given by:
@@
2.5.3 Quartic Smoothing Spline Model
A non-parametric framework called the Quartic Smoothing Spline (QSS) was created to recover the intrinsic temporal structure of GRB LCs from observational data that is noisy and irregularly sampled. A piecewise fourth-degree polynomial basis, which offers a potent blend of high-order smoothness and local flexibility, defines the model’s architecture. Instead of minimizing a single penalized functional, the model finds the optimal representation of the LC, $S(x)$, by solving a constrained optimization problem. It minimizes the sum of squared residuals to the data, subject to a smoothness constraint based on the spline’s derivatives (1993csfw.book.....D):
$\min_{S}\sum_{i=1}^{N}\bigl[y_{i}-S(x_{i})\bigr]^{2}\quad\text{subject to}\quad\sum_{j}\left(\Delta^{(5)}S(\tau_{j})\right)^{2}\leq s,$ (24)
where $(x_{i},y_{i})$ represent the observed log-time and log-flux data pairs. The novelty of this framework lies in its definition of smoothness: for a quartic spline ($k=4$), the constraint is placed on the sum of the squared discontinuity jumps of the fifth derivative ($\Delta^{(5)}S$) at each interior knot $\tau_{j}$. This provides a powerful modeling dynamic where the spline’s local flexibility is regularized by a high-order smoothness condition, a trade-off well-suited for the complex morphology of GRBs.
The construction and implementation of the QSS model are governed by the following structural choices and procedures:
  • 1.
    Basis Representation: The fundamental building block of the model is a basis of degree-four ($k=4$) B-spline functions (deboor2001practical). The final spline is a linear combination of these basis functions:
    $S(x)\;=\;\sum_{j=1}^{n}c_{j}\,B_{j}(x),$ (25)
    where $\{B_{j}(x)\}$ are the quartic B-spline basis functions and the coefficients $\{c_{j}\}$ are determined via the constrained optimization. This architectural choice endows the model with its characteristic properties.
  • 2.
    Inherent Structural Properties: The degree-four basis offers two significant features. The first is enhanced local flexibility, which allows the model to accurately represent complex morphological features such as sharp, asymmetric peaks and flat plateaus. Because the resulting curve is three times continuously differentiable ($C^{3}$), the second is higher-order smoothness, which ensures that the rate of change of curvature is continuous and can represent more physically plausible transitions in the GRB emission process.
  • 3.
    Smoothing and Regularization: The smoothing factor, $s$, serves as the upper bound for the smoothness constraint in the optimization problem. With $N$ representing the number of data points, we set $s=N$. For astrophysical time-series data, this choice offers an empirical balance between bias and variance, making it a robust method for regularization (hastie2009elements).
  • 4.
    Implementation and Evaluation: The model is implemented using the scipy.interpolate.UnivariateSpline routine. The solver places knots at the data locations and efficiently computes the coefficient vector $\mathbf{c}$. The final, continuous representation $S(x)$ is then evaluated on a dense grid to produce the high-resolution reconstructed LC.
To provide a complete probabilistic reconstruction, we model the uncertainty inherent in the observations. The residuals between the fitted spline and the data, $r_{i}=y_{i}-S(x_{i})$, are calculated to estimate a residual standard deviation, $\sigma_{\rm resid}$:
$\sigma_{\rm resid}\;=\;\sqrt{\frac{\sum_{i=1}^{N}r_{i}^{2}}{N}}.$
To generate realistic synthetic data points, we first characterize the original observational log-flux errors by fitting both Normal and Laplace distributions and selecting the one with the highest likelihood. Noise is then sampled from this best-fit distribution to be added to the mean spline prediction. For uncertainty visualization, 95% confidence bands are constructed around the mean spline fit as $S(x)\;\pm\;1.96\,\sigma_{\rm resid}$, which are then transformed from log-space back to linear flux units for physical interpretation (eilers1996flexible). The model’s generalization performance is validated using 5-fold cross-validation. The entire workflow is summarized in Fig. 8.
Load Time & Flux DataLog–Log TransformNormalize to Zero Mean & Unit VarianceFit Quartic Spline (k=4, s=N)Evaluate on Dense GridCompute Residuals & $\sigma_{\rm resid}$Sample & Add NoiseConstruct 95% CIOutput Reconstructed Curve
Figure 8: Workflow of the quartic smoothing spline model for GRB light‐curve reconstruction, from raw data loading through spline fitting to uncertainty quantification.
2.6 Parameter Uncertainty and Evaluation Metrics
In training our reconstruction model, we do not impose any analytical form on the GRB afterglow, such as the W07 profile or a broken power law. The machine learning framework is therefore inherently model-independent, allowing it to learn the temporal structure directly from the data rather than being constrained by a predetermined shape. Nevertheless, to determine whether the reconstruction leads to improved physical parameter estimates, a quantitative benchmark is required.
For this purpose, we use the W07 as a reference and take its fitted parameters as the baseline for comparison. Following 2020ApJ...903...18S, the uncertainty associated with each parameter $p=(\log T_{a},\log F_{a},\alpha)$, are defined as:
$\Delta p=\frac{\max(p)-\min(p)}{2}.$ (26)
It is clear that with data augmentation, since the uncertainties, $\delta$ scales as $1/\sqrt{N}$ where N is the number of data points, in principle one can increase this number, and thus reduce the uncertainty. However, we have been very careful in data augmentation, because we aim to preserve the features of the initial LCs so that the increase in the number of data points should be proportional to the existing data points. Indeed, if we increase the number of data points without accounting for this recipe, systematic biases will result. Thus, we have worked out a strategy that balances data augmentation, preserves the original model we assume (in this case, the (willingale2007testing) function) when the LCs are reconstructed, and still reduces the final uncertainties in the given model’s parameters. For details about the method used, see Appendix B. This configuration may represent current data only if we assume the satellite we are considering does not undergo any orbital gap. So, data augmentation will solve the issue of gaps if perfect seeing will be available. One can also consider the same situation if you allow future satellites that, when combined, provide full data coverage. This analysis was performed using the W07 model, a benchmark. Indeed, the same procedure can be repeated for any other model, as we have already done in (dainotti2023stochastic), in which a broken power-law was tested. The additional model testing or simulations of future satellites are beyond the scope of the current paper.
This definition captures the observational limitations inherent in GRB afterglow measurements, where incomplete LC coverage introduces non-negligible statistical uncertainty. In the absence of fully sampled LCs, this metric provides a consistent basis for evaluating whether our reconstruction effectively reduces parameter uncertainties relative to those obtained from the W07 model.
To quantify this improvement, we compute the error fraction (EF) for each parameter before and after reconstruction dainotti2023stochastic: